M.V.P. Games, Inc., has hired you to perform a feasibility study of a new video game that requires an initial investment of $5,676,800. The company expects a total annual operating cash flow of $1.22 million for the next 9 years. The relevant discount rate is 10 percent. Cash flows occur at year-end. After one year, the estimate of remaining annual cash flows will be revised either upward to $2.12 million with a probability of 30% or downward to $277,000 with a probability of 70%. At that time, the video game project can be sold for $2.52 million. Should the firm abandon the project after one year? What is the revised NPV?

Yes. 995,474.34

Yes. 3,136,864.07

No.1,477,774.56

No. 626,009.06
Answer #1

Answer :

Initial Investment i = - 5,676,800

Life of Project n = 9 years

CF1 = $1.22 million

Discount Rate   r = 0.10

CF2 = 2,120,000*0.3 + 277,000*0.7 = 636000+193900= 829,900

At the end of 1 year, accumulated benefit from the project => Present Value of Cash flows

=> CF1/(1+r)1 = 1220000/(1.10)1 = 1,109,090.9

Part of initial investment recovered from cash flows in 2 years= 1109091

Cost to be recovered= 5676800-1109091 = 4567709

Net Loss= Cost Price- Selling Price= 4567709 - ed432520000 = 2047709

Comparing this with possible loss/profit if the project is continued.

Present Value Annuity Factor of Ordinary Annuity at 10% for 9 years= 5.75902

Present Value Annuity Factor of Ordinary Annuity at 10% for 1 year= 0.90909

Present Value Annuity Factor of Ordinary Annuity at 10% for the period from 2nd year to 9th year = PFAF9 - PFAF1

Net PFAF = 5.75902 - 0.90909 = 4.84993

Discounted Cash flows (Revised) for 8 years = 829900*4.84993 = 4024956.91 ~ 4,024,957

Discounted Cash flow for the 1st year = 1109090

Total Discounted Cash flows = 5134047.81

NPV= DCF9 - i = 5134048-5676800 = - 542750

No, the project should not be abandoned.